# In Search of Adam’s Secret Sauce

### v1

@@ Abstract
Understanding the remarkable efficacy of Adam when training transformer-based language models has become a central research topic within the optimization community. To gain deeper insights, several simplifications of Adam have been proposed, such as the signed gradient and signed momentum methods. In this work, we conduct an extensive empirical study — training over 1,300 language models across different data configurations and scales — comparing Adam to several known simplified variants. We find that signed momentum methods are faster than SGD, but consistently underperform relative to Adam, even after careful tuning of momentum, clipping setting and learning rates. However, our analysis reveals a compelling option that preserves near-optimal performance while allowing for new insightful reformulations: constraining the Adam momentum parameters to be equal, $\beta_{1}=\beta_{2}$. Beyond robust performance, this choice affords new theoretical insights, highlights the “secret sauce” on top of signed momentum, and grants a precise statistical interpretation: we show that Adam in this setting implements a natural online algorithm for estimating the mean and variance of gradients—one that arises from a mean-field Gaussian variational inference perspective.
@@ 1 Introduction
Table 1:  (Signum closes 96% of the perplexity gap between Adam and SGD) Validation perplexity comparison of widely used optimizers that interpolate between SGD and Adam, evaluated on a language modeling task (160M parameters, 3.2B SlimPajama tokens, sequence length 2048, batch size 256 – Chinchilla optimal). We report the mean and 2-sigma interval of validation perplexity (on 100M held-out tokens) across 3 initialization seeds. Weight decay is always decoupled  [Loshchilov and Hutter, 2019] and set to $0.1$ [Biderman et al., 2023, Liu et al., 2024] except for SGD where we further tune (§B). RMSprop does not use momentum, and Gclip is global norm clipping to $1$ (before applying momentum), Cclip is coordinate-wise clipping (after applying momentum). Other hyperparameters, for all other methods, are carefully tuned, see e.g. Figure 2 and §3.
@@
Val ppl. 21.86$\pm$ 0.21 23.23$\pm$ 0.16 27.04$\pm$ 0.34 33.40$\pm$ 0.39 36.78$\pm$ 0.57 37.76$\pm$ 0.61 53.62$\pm$ 5.14
While for large-scale training, the slowdown in Figure 1 is not acceptable, it may seem unnecessary or anachronistic to further explain it, in light of recent algorithms claiming to have further improved the performance of Adam, e.g. Muon [Jordan et al., 2024, Liu et al., 2025, Shah et al., 2025], Scion [Pethick et al., 2025], and Shampoo-based [Gupta et al., 2018] methods such as SOAP [Vyas et al., 2025]. However, a close inspection of such optimizers reveals that, while gains over vanilla Adam are solid, those are contingent on the use/adaptation of Adam on a specific subset of parameters: For instance, in recent scaled-up versions of Muon [Liu et al., 2025, Shah et al., 2025], Adam is used to update embedding, LM heads and normalization parameters  [^1] Coincidentally, the ones that were shown to be most sensitive during training [Zhao et al., 2025, Kunstner et al., 2024]. Scion, claiming a greater independence from Adam, adopts an architecture where normalization layers have no gain parameters., and on the other parameters the Muon update is normalized to have a similar RMS value similar to the Adam update. Further, SOAP’s improvements stem from the application of Adam in the preconditioner’s eigenbasis.
@@
Figure 2: Training a total of 265 language models with 160M parameters with 3.2B SlimPajama-627B tokens, sequence length of 2048, batch size of 256. Shown is the final test perplexity on 100M held-out tokens. Some underperforming runs are not shown to keep focus on the most interesting range. For a careful description of our tuning grid, see §A. Takeaway 1: Validation perplexity of highly tuned (65 hyperparameter configurations) Signum with weight decay 0.1 – top row – is around 23.23 (see Table 1 for multiple seeds at optimal tuning). We ablate on the momentum parameter, learning rate, and presence of global clipping before averaging. The best performance of Signum is reported as a green horizontal line on the second row (200 Adam runs, with weight decay of $0.1$). Most Adam runs perform better than optimally tuned Signum. Takeaway 2: For each $\beta_{1}$, the optimal corresponding $\beta_{2}$ (after learning rate tuning) is similar. The higher $\beta_{1}$, the higher $\beta_{2}$ for optimal performance (optimal $\beta$s are correlated).
@@ 2 Preliminaries and Related Works
[…]
@@
$\texttt{clip}[g_{k}]=\min\left\{1,\frac{1}{\|g_{k}\|_{2}}\right\}g_{k}.$
@@ 3.1 Extensive benchmarking at 160M parameters
To put to the test our second takeaway in different training settings, we consider shorter sequence lengths (512, Figure 12), higher/lower batch sizes (Figure 14 & Figure 15), different data (Fineweb) and absence of weight decay (Fig, 16). See discussion in §3.2.
@@ 3.2 Ablations
Different batch sizes.
@@
We find our Takeaway 2 to also hold at shorter sequence length of 512 (Figure 12). We note that here performance of Signum is closer to that of Adam compared to Figure 2 – yet, Adam is still superior by a substantial margin ( 0.7 validation perplexity), Takeaway 1. This pattern agrees well with the results in [Zhao et al., 2025], who found other methods to be competitive with Adam at short context lengths. Our experiments in Figure 2 and Figure 16 suggest that Adam performance particularly shines at higher sequence lengths.
@@
Larger Models
We restrict out attention to the SlimPajama dataset and to validation of Takeaway 2. Results are presented in Figure 4, comprising 44 full compute-optimal training runs of 410M parameter models, which confirm yet again strong and near-optimal performance at $\beta_{1}=\beta_{2}$.
Figure 4: The final validation performance (100M held-out tokens) for 44 trained LMs with 420M parameters trained on 8.2 B SlimPajama tokens (Chinchilla-optimal). Equal betas yields near-optimal performance. We use gradient clipping and a batch size of 512 (scaled by 2 compared to Figure 2, as suggested by Zhang et al. [2025]). Sequence length is 2048, weight decay is $0.1$. Note that the standard setting $(0.9,0.95)$ is quite suboptimal here.
@@ 4 New Viewpoints of Adam
We next rewrite the update in a form that explicitly highlights the role of variance.
@@
Comparison with Balles and Hennig [2018].
Setting $\beta_{1}=\beta_{2}$, a choice that was not common [^4] Default parameters have for long been $\beta_{1}=0.9$, $\beta_{2}=0.999$, see https://docs.pytorch.org/docs/stable/generated/torch.optim.Adam.html. at the time of Balles and Hennig [2018] allows for precise calculations and claims around variance estimation. In particular, the quantity $\texttt{EMA}_{\beta_{2}}[g_{k}^{2}]-\texttt{EMA}_{\beta_{1}}[g_{k}]^{2}$ allows for a precise variance interpretation only for the case $\beta_{1}=\beta_{2}$: we prove in §C.2 that for any $a,b,\gamma\in\mathbb{R}$ and $\tau\in(0,1)$, the representation
$d_{k}=\frac{m_{k}}{\sqrt{m_{k}^{2}+\gamma\,\texttt{EMA}_{\tau}[(am_{k-1}-bg_{k})^{2}]}}$
holds for Adam if and only if $\beta_{1}=\beta_{2}$. For a proof, see §C.2 in the appendix. In other words, this shows that linking Adam to variance estimation can only be done precisely for the case of equal betas.
@@ 4.1 Adam Estimates Mean and Variance using Variational Inference
We now show that Adam admits a natural interpretation as an online variational inference method, in which
$m_{k}:=\texttt{EMA}_{\beta}[g_{k}]\quad\text{and}\quad\sigma_{k}:=\beta\,\texttt{EMA}_{\beta}[(m_{k-1}-g_{k})^{2}]$
@@
$\min_{m,\sigma^{2}\geq 0}\;-\log p(g_{k+1}\mid m,\sigma^{2})+\lambda\,\mathrm{KL}\left(\mathcal{N}(m_{k},\sigma^{2}_{k})\,\|\,\mathcal{N}(m,\sigma^{2})\right),$ (4)
@@
The following result characterizes the solution of (4), showing that the moving averages used in Adam correspond exactly to an instance of online variational inference:
@@
$m_{k+1}$ $=\beta m_{k}+(1-\beta)g_{k+1}=\texttt{EMA}_{\beta}(g_{k+1}),$ (7)
$\sigma^{2}_{k+1}$ $=\beta\sigma^{2}_{k}+\beta(1-\beta)(m_{k}-g_{k+1})^{2}=\beta\,\texttt{EMA}_{\beta}\left((m_{k}-g_{k+1})^{2}\right).$ (8)
@@
$d_{k}\;=\;\frac{\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}}{\sqrt{{\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}^{2}}+{\color[rgb]{0.62890625,0.234375,0.203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0.234375,0.203125}\beta\texttt{EMA}{{}_{\beta}}((m_{k-1}-g_{k})^{2})}}}\;=\;\frac{\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}}{\sqrt{{\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}^{2}}+{\color[rgb]{0.62890625,0.234375,0.203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0.234375,0.203125}\sigma^{2}_{k}}}}=\frac{\operatorname{sign}({\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}})}{\sqrt{1+{\color[rgb]{0.62890625,0.234375,0.203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0.234375,0.203125}\sigma_{k}^{2}}/{\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}^{2}}}}.$ (9)
@@
Using these insights, we can better formalize the noise-to-signal interpretation of Adam [Balles and Hennig, 2018]. Let the quantity $m_{k}/\sigma_{k}$ denote the signal-to-noise ratio (SNR). We show that Adam can be viewed as a steepest descent method whose trust region is modulated by the SNR.
@@
$-\mbox{sign}(m_{k})\;=\;\underset{\theta\in\mathbb{R}}{\operatorname{argmin}}\;-m_{k}\cdot\theta\quad\text{subject to }|\theta|\leq 1.$ (10)
That is, Signum selects the direction most aligned with $-m_{k}$ within a unit trust region.
@@
$-\frac{\mbox{sign}(m_{k})}{\sqrt{1+\sigma_{k}^{2}/m_{k}^{2}}}\;=\;\underset{\theta\in\mathbb{R}}{\operatorname{argmin}}\;-m_{k}\cdot\theta\quad\text{subject to }|\theta|\leq\frac{1}{\sqrt{1+\sigma_{k}^{2}/m_{k}^{2}}}.$ (11)

### v2

@@ Abstract
Understanding the remarkable efficacy of Adam when training transformer-based language models has become a central research topic within the optimization community. To gain deeper insights, several simplifications of Adam have been proposed, such as the signed gradient and signed momentum methods. In this work, we conduct an extensive empirical study — training over 1,500 language models across different data configurations and scales — comparing Adam to several known simplified variants. We find that signed momentum methods are faster than SGD, but consistently underperform relative to Adam, even after careful tuning of momentum, clipping setting and learning rates. However, our analysis reveals a compelling option that preserves near-optimal performance while allowing for new insightful reformulations: constraining the Adam momentum parameters to be equal, $\beta_{1}=\beta_{2}$. Beyond robust performance, this choice affords new theoretical insights, highlights the secret sauce on top of signed momentum, and grants a precise statistical interpretation: we show that Adam in this setting implements a natural online algorithm for estimating the mean and variance of gradients—one that arises from a mean-field Gaussian variational inference perspective.
@@ 1 Introduction
Table 1:  (Signum closes 96% of the perplexity gap between Adam and SGD) Validation perplexity comparison of widely used optimizers that interpolate between SGD and Adam, evaluated on a language modeling task (160M parameters, 3.2B SlimPajama tokens, sequence length 2048, batch size 256 – Chinchilla optimal). We report the mean and 2-sigma interval of validation perplexity (on 100M held-out tokens) across 3 initialization seeds. Weight decay is always decoupled  [Loshchilov and Hutter, 2019] and set to $0.1$ [Biderman et al., 2023, Liu et al., 2024] except for SGD where we further tune (§B). RMSprop does not use momentum, and Gclip is global norm clipping to 11 (before applying momentum), Cclip is coordinate-wise clipping (after applying momentum). Other hyperparameters, for all other methods, are carefully tuned, see e.g. Figure 2 and §3.
@@
[…]
While for large-scale training, the slowdown in Figure 1 is not acceptable, it may seem unnecessary or anachronistic to further explain it, in light of recent algorithms claiming to have further improved the performance of Adam, e.g. Muon [Jordan et al., 2024, Liu et al., 2025, Shah et al., 2025], Scion [Pethick et al., 2025], and Shampoo-based [Gupta et al., 2018] methods such as SOAP [Vyas et al., 2025]. However, a close inspection of such optimizers reveals that, while gains over vanilla Adam are solid, most of these methods still use Adam on a specific subset of parameters: For instance, in recent scaled-up versions of Muon [Liu et al., 2025, Shah et al., 2025], Adam is used to update embedding, LM heads and normalization parameters  [^1] Coincidentally, the ones that were shown to be most sensitive during training [Zhao et al., 2025, Kunstner et al., 2024]. Scion claims a greater independence from Adam, yet adopts an architecture where normalization layers have no trainable gain parameters. While results are promising, experiments in the usual setup are needed., and on the other parameters the Muon update is normalized to have a similar RMS value similar to the Adam update. Further, SOAP’s improvements stem from the application of Adam in the preconditioner’s eigenbasis.
@@
[…]
@@ 2 Preliminaries and Related Works
$\texttt{EMA}_{\beta}[s_{k}]=\beta\texttt{EMA}_{\beta}[s_{k-1}]+(1-\beta)s_{k},\qquad\texttt{EMA}_{\beta}[s_{0}]:=s_{0}\ \text{(or zero)}.$ (1)
@@
$\texttt{Gclip}[g_{k}]=\min\left\{1,\frac{1}{\|g_{k}\|_{2}}\right\}g_{k}.$
@@ 3.1 Extensive benchmarking at 160M parameters
To put to the test our second takeaway in different training settings, we consider shorter sequence lengths (512, Figure 14), higher/lower batch sizes (Figure 16 & Figure 17), different data (Fineweb) and absence of weight decay (Fig, 18). See discussion in §3.2.
@@
Theoretical relations between betas. We note that a correlation between $\beta$ parameters was also noted first by Reddi et al. [2018], Alacaoglu et al. [2020] for AMSgrad, and later by Zhang et al. [2022] for Adam, where it is shown that if $\beta_{2}$ is large enough and $\beta_{1}<\sqrt{\beta_{2}}$, it converges to the neighborhood of critical points. Further, Xie and Li [2024] showed that weight decay in AdamW leads to convergence to a constrained minimizer only if $\beta_{2}>\beta_{1}$.
@@ 3.2 Ablations
More Tokens.
We find our Takeaway 2 to also hold at a higher token budget. In §B.2, we show a trend very similar to Fig. 2 for models trained for $2\times$ the Chinchilla-optimal budget.
Different batch size.
@@
In §B.3, we find our Takeaway 2 to also hold at shorter sequence length of 512 (Figure 14). We note that here performance of Signum is closer to that of Adam compared to Figure 2 – yet, Adam is still superior by a substantial margin ( 0.7 validation perplexity), Takeaway 1. This pattern agrees well with the results in [Zhao et al., 2025], who found other methods to be competitive with Adam at short context lengths. Our experiments in Figure 14 and Figure 2 suggest that Adam performance particularly shines at higher sequence lengths.
@@
Larger Models.
We restrict our attention to the SlimPajama dataset and to validation of Takeaway 2. Results are presented in Figure 5, comprising 44 full compute-optimal training runs of 410M parameter models, which confirm yet again strong and near-optimal performance at $\beta_{1}=\beta_{2}$.
Figure 5: The final validation performance (100M held-out tokens) for 44 trained LMs with 410M parameters trained on 8.2 B SlimPajama tokens (Chinchilla-optimal). Equal betas yields near-optimal performance. We use gradient clipping and a batch size of 512 (scaled by 2 compared to Figure 2, as suggested by Zhang et al. [2025]). Sequence length is 2048, weight decay is $0.1$. Note that the standard setting $(0.9,0.95)$ is quite suboptimal here.
@@ 4 New Viewpoints of Adam
We next rewrite (proof in the Appendix) the update to explicitly highlight the role of variance.
@@ 4.1 Adam Estimates Mean and Variance using Variational Inference
We show that Adam admits a natural interpretation as an online variational inference method, where
$m_{k}:=\texttt{EMA}_{\beta}[g_{k}]\quad\text{and}\quad\sigma_{k}^{2}:=\beta\,\texttt{EMA}_{\beta}[(m_{k-1}-g_{k})^{2}]$
@@
$\min_{m,\sigma^{2}\geq 0}\;-\log p(g_{k+1}\mid m,\sigma^{2})+\,\tfrac{1}{\lambda}\mathrm{KL}\left(\mathcal{N}(m_{k},\sigma^{2}_{k})\,\|\,\mathcal{N}(m,\sigma^{2})\right),$ (4)
@@
The following result, proved in the appendix, characterizes the solution of (4), showing that the moving averages used in Adam correspond exactly to an instance of online variational inference:
@@
$m_{k+1}$ $=\beta m_{k}+(1-\beta)g_{k+1}=\texttt{EMA}_{\beta}[g_{k+1}],$ (7)
$\sigma^{2}_{k+1}$ $=\beta\sigma^{2}_{k}+\beta(1-\beta)(m_{k}-g_{k+1})^{2}=\beta\,\texttt{EMA}_{\beta}\left[(m_{k}-g_{k+1})^{2}\right].$ (8)
@@
$d_{k}\;=\;\frac{\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}}{\sqrt{{\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}^{2}}+{\color[rgb]{0.62890625,0.234375,0.203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0.234375,0.203125}\beta\texttt{EMA}{{}_{\beta}}[(m_{k-1}-g_{k})^{2}]}}}\;=\;\frac{\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}}{\sqrt{{\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}^{2}}+{\color[rgb]{0.62890625,0.234375,0.203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0.234375,0.203125}\sigma^{2}_{k}}}}=\frac{\operatorname{sign}({\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}})}{\sqrt{1+{\color[rgb]{0.62890625,0.234375,0.203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0.234375,0.203125}\sigma_{k}^{2}}/{\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}^{2}}}}.$ (9)
@@
Using these insights, we can better formalize the noise-to-signal interpretation of Adam [Balles and Hennig, 2018] (see also §4.2). Let $m_{k}/\sigma_{k}$ denote the signal-to-noise ratio (SNR). We show that Adam can be viewed as a steepest descent method whose trust region is modulated by the SNR.
@@
$\mbox{sign}(m_{k})\;=\;\underset{\theta\in\mathbb{R}}{\operatorname{argmin}}\;-m_{k}\cdot\theta\quad\text{subject to }|\theta|\leq 1.$ (10)
That is, Signum selects the direction most aligned with $m_{k}$ within a unit trust region.
@@
[…]
@@
4.2 Comparison with Balles and Hennig [2018]
Balles and Hennig [2018] first drew a connection between Adam, Signum and Signal-to-noise Ratio regularization. Their observation was as follows. Let $m_{k}=\texttt{EMA}_{\beta_{1}}[g_{k}]$, and $v_{k}=\texttt{EMA}_{\beta_{2}}[g_{k}^{2}]$. We can trivially re-write the Adam direction as
$d_{k}=\frac{m_{k}}{\sqrt{v_{k}}}={\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}\frac{m_{k}}{\sqrt{m_{k}^{2}+{\color[rgb]{0.62890625,0.234375,0.203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.62890625,0.234375,0.203125}v_{k}-m_{k}^{2}}}}}.$
If we now assume that $\sigma_{k}^{2}:=v_{k}-m_{k}^{2}$ is a measure of variance, then dividing the Adam direction through by $|{\color[rgb]{0.1328125,0.46875,0.33203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.46875,0.33203125}m_{k}}|$, as done in (9), we arrive at a Signal-to-noise Ratio regularized variant of the Signum method. In particular, as the noise goes to zero ($\sigma_{k}^{2}\rightarrow 0$), we arrive at the Signum method.
The missing piece in their insight was to show when and if the term $v_{k}-m_{k}^{2}$ is a measure of variance.
We show that $\beta_{1}=\beta_{2}$, a choice that was not common [^4] Default parameters have for long been $\beta_{1}=0.9$, $\beta_{2}=0.999$, see https://docs.pytorch.org/docs/stable/generated/torch.optim.Adam.html. at the time of Balles and Hennig [2018], allows for more precise claims: Proposition 1 shows that when $\beta_{1}=\beta_{2}=\beta$ the term $v_{k}-m_{k}^{2}$ is precisely equal to $\beta\texttt{EMA}_{\beta}[(m_{k-1}-g_{k})^{2}]$, which in turn is a online estimate of variance (Theorem 4.1). We further show that $v_{k}-m_{k}^{2}$ only has a precise variance interpretation for the case $\beta_{1}=\beta_{2}$: indeed, we prove in §C.2 that Adam can be represented as
$d_{k}=\frac{m_{k}}{\sqrt{m_{k}^{2}+\gamma\,\texttt{EMA}_{\tau}[(am_{k-1}-bg_{k})^{2}]}}$ (12)
for some $a,b,\gamma\in\mathbb{R}$ and $\tau\in(0,1)$ if and only if $\beta_{1}=\beta_{2}$. In other words, connecting $v_{k}-m_{k}^{2}$ to variance estimation, and in turn Adam to an SNR-controlled trust region method (11), can only be done precisely for the case of equal betas.
Ablating hyperparameters in our reformulation.
While (12) reduces to Adam with equal betas if and only if $a,b=1$ and $\beta=\gamma=\tau$, we found it interesting to consider (12), with $a=b=1$, as a new method with no precise connection to simultaneous variance and mean estimation, with hyperparameters $\beta,\gamma,\tau$. In §C.4, we train 150 additional language models ablating on such parameters, and found no advantage in setting $\beta\neq\tau$ or $\tau\neq\gamma$. We believe such evidence further strengthens our claims: best performance is aligned to the theoretical choice $\tau=\gamma=\beta$.